# Improving the Language Understanding Capabilities of Large Language Models Using Reinforcement Learning

### v4

@@ Abstract
Instruction-fine-tuned large language models (LLMs) under 14B parameters continue to underperform on natural language understanding (NLU) tasks, often trailing smaller models like BERT-base on benchmarks such as GLUE and SuperGLUE. Motivated by the success of reinforcement learning in reasoning tasks (e.g., DeepSeek), we explore Proximal Policy Optimization (PPO) as a framework to improve the NLU capabilities of LLMs. We frame NLU as a reinforcement learning environment, treating token generation as a sequence of actions and optimizing for reward signals based on alignment with ground-truth labels. PPO consistently outperforms supervised fine-tuning, yielding an average improvement of 6.3 points on GLUE, and surpasses zero-shot and few-shot prompting by 38.7 and 26.1 points, respectively. Notably, PPO-tuned models outperform GPT-4o by over 4% on average across sentiment and natural language inference tasks, including gains of 7.3% on the Mental Health dataset and 10.9% on SIGA-nli. This work highlights a promising direction for adapting LLMs to new tasks by reframing them as reinforcement learning problems, enabling learning through simple end-task rewards rather than extensive data curation.
@@
Sai Ashish Somayajulafootnotemark:  Bokai Hu [^1] Equal Contribution  Xin Pan  Pengtao Xie UC San Diego, USA {ssomayaj, p1xie}@ucsd.edu
@@ 1 Introduction
Figure 1: PPO-based fine-tuning of LLAMA2-7B-chat-hf to improve the performance on NLU tasks.
Large language models (LLMs) (Radford et al., 2019; Brown, 2020; Touvron et al., 2023b) have revolutionized natural language processing (NLP) with their powerful text generation capabilities (Radford, 2018). Pretrained on large-scale unlabeled corpora, LLMs can generate coherent and contextually relevant content. Using prompt-based strategies like zero-shot and few-shot prompting (Brown, 2020), these models can address a wide range of downstream tasks without task-specific fine-tuning. However, when applied to instruction-fine-tuned LLMs under 14B parameters—such as LLAMA2-7B-chat-hf—these methods often underperform on natural language understanding (NLU) tasks compared to encoder-only models like BERT (Devlin, 2018), which consistently excel on benchmarks such as GLUE (Wang et al., 2019) and SuperGLUE (Wang et al., 2020). For instance, our evaluation of LLAMA2-7B-chat-hf shows that zero-shot prompting with task-specific prompts yields an average performance of 46.1 across all GLUE datasets, while few-shot prompting improves performance to 58.7—both significantly trailing BERT-base’s 79.6, as shown in Table 1.
@@
While standard fine-tuning (SFT) is commonly used to adapt LLMs to downstream tasks, we find it insufficient for NLU—often underperforming even smaller encoder-only models like BERT-base. In contrast, we use PPO to enhance LLM performance by framing NLU as a reinforcement learning problem. The sequence of input tokens up to timestep $t-1$ represents the state $s_{t}$, and the token generated at timestep $t$ is treated as the action $a_{t}$. After generating the full response, a heuristic extracts the predicted answer, which is compared to the ground-truth label to assign a scalar reward $R$. PPO then updates the model to maximize this reward, enabling direct optimization for task-specific objectives. Empirically, PPO-based fine-tuning of LLAMA2-7B-chat-hf improves GLUE performance by 6.3 points over SFT, surpasses zero- and few-shot prompting by 38.7 and 26.1 points, and even outperforms GPT-4o by over 4% on sentiment and inference tasks—achieving gains of 7.3% on the Mental Health dataset and 10.8% on SIGA-nli. These results demonstrate the effectiveness of reinforcement learning in aligning LLMs under 14B parameters with NLU objectives.
@@
Pre-trained LLMs possess broad linguistic knowledge—spanning syntactic and semantic structures—acquired from large-scale text corpora. We show that reinforcement learning, specifically PPO, can refine this general understanding to better align with task-specific NLU objectives. Similar patterns are observed in DeepSeek, where chain-of-thought pre-training enhances reasoning capabilities, and subsequent RL-based fine-tuning further improves performance. Building on this insight, our findings suggest a promising direction: adapting LLMs to new tasks by formulating them as reinforcement learning problems. When models are sufficiently pre-trained, task alignment may be achieved without additional labeled data—requiring only a well-defined reward function over the outputs. PPO can then optimize the model toward high-reward behaviors. This approach offers a scalable, label-efficient alternative to conventional supervised fine-tuning through reward-driven adaptation.
2 Related Works
3 Preliminaries on Application of PPO for Fine-tuning LLMs
Proximal policy optimization (PPO)(Schulman et al., 2017b) is an online reinforcement learning algorithm. In this section, we describe the process to fine-tune an LLM using PPO. During training, at each timestep $t$, the LLM (policy) generates a token prediction $a_{t}$ (action) based on the state $s_{t}$, which consists of the sequence of generated tokens up to timestep $t-1$. The final generated output is evaluated in the context of the downstream task, where the environment provides feedback in the form of rewards. The model updates its parameters based on these rewards to improve its ability to generate accurate predictions over time.
[…]
Here, $p_{\theta}(a_{t}|s_{t})$ is the probability of taking action $a_{t}$ in state $s_{t}$ under the current policy, while $p_{\theta^{\prime}}(a_{t}|s_{t})$ represents this probability under the old policy. In PPO, the training data—specifically, the state-action pairs $(s_{t},a_{t})$—are sampled using the old policy $\pi_{\theta^{\prime}}$ (the LLM before it is updated), rather than the new policy currently being optimized. Thus, the ratio $\frac{p_{\theta}(a_{t}|s_{t})}{p_{\theta_{\text{old}}}(a_{t}|s_{t})}$ accounts for how much the new policy has changed relative to the old policy and adjusts the likelihood of an action accordingly. This ratio is multiplied by $\hat{A}_{t}$, the Generalized Advantage Estimation (GAE) (Schulman et al., 2018), which measures how much better or worse an action $a_{t}$ is compared to the expected outcome under the current policy.
$\hat{A}_{t}=R_{t}+\gamma V_{t+1}-V_{t}+\gamma\lambda\hat{A}_{t+1},$
Critic Model
$\delta_{t}=(R_{t}+\gamma V_{t+1}-V_{t})^{2},$ (2)
@@ 4 Method
To enhance the performance of LLMs on NLU tasks, we adopt two distinct fine-tuning methods. The first approach involves supervised fine-tuning, where the input consists of a concatenation of the task-specific prompt, query and the ground truth answer, with the model optimized using the next-token prediction objective. The second approach utilizes PPO, framing response generation as a reinforcement learning problem. In this setup, the sequence of input tokens until timestep $t-1$ represents the state $s_{t}$, and each token generated at timestep $t$ is treated as an action $a_{t}$. After generating the entire sequence, a heuristic-based process extracts the final answer from this generated sequence, and is compared to the ground truth. PPO is then employed to optimize the model by maximizing the cumulative reward derived from this comparison. To reduce computational complexity, we fine-tune LoRA layers instead of the full model.
@@ 4.3 Proximal Policy Optimization for LLM Fine-tuning on NLU Tasks
We utilize PPO to fine-tune the LLM on NLU tasks, following the training protocol outlined in section 3. The reward function is specifically designed for each NLU task. In this work, we use a simple reward function, where a reward is assigned at the end of the generation based on alignment with the ground truth labels. We use regular expression matching to extract answers from the LLMs outputs by first locating the text within the ‘<Judgement></Judgement>’ tags. Depending on the task, we then search for task-specific keywords (such as “yes”, “no”, “acceptable”, or “not acceptable”) to identify the answer. These extracted answers are compared with the ground truth to determine the appropriate rewards.
For instance, CoLA, which is a classification task, answers are categorized as acceptable, unacceptable, or exceptional (incorrect format). For STS-B, which is a regression task, the extracted answer is a floating-point number between 0 and 5. The reward per generation for classification tasks is given by $R=\mathds{1}(\hat{y}==y_{i})$, where $\hat{y}$ is the model’s prediction and $y$ is the ground truth. For STS-B, a regression task, the reward per generation is calculated based on how close the prediction is to the ground truth: $R=2.5-|\hat{y_{i}}-y_{i}|$. Incorrectly formatted responses are penalized with a value of -1 for classification tasks and -2.5 for regression tasks.
@@ 5.1 Experimental Setup
We trained and evaluated our models on the GLUE(Wang et al., 2019) and SuperGLUE(Wang et al., 2020) benchmarks. All experiments were conducted using instruction-tuned LLAMA2-7B models(Touvron et al., 2023a) [^1] https://huggingface.co/daryl149/llama-2-7b-chat-hf. We perform both single task and multi-task fine-tuning: 1) Single-task Fine-tuning: For each subtask within GLUE and SuperGLUE, a separate task-specific LoRA module was trained independently. 2) Multi-task Fine-tuning: In the multi-task setting, datasets from different subtasks within each benchmark were combined, and a single LoRA module was trained to handle all tasks simultaneously. Please refer to Appendix A for detailed hyperparameter settings.
@@ 5.3 Results on GLUE Benchmark
Models MNLI-m MNLI-mm QQP QNLI SST-2 CoLA
@@ 5.4 Evaluating Zero-Shot Generalization of PPO Fine-Tuned Models and Comparison with GPT-4o
Tasks LLAMA2-7B PPO-ST LLAMA2-13B PPO-ST GPT-4o
Sentiment Analysis
Financial PhraseBank 97.2 97.7 97.5
Labelled Financial News 70.2 72.3 67.8
Mental Health 67.2 66.6 59.9
Emotion 78.0 76.4 77.6
Natural Language Inference
Babi-nli 68.3 69.4 63.2
SIGA-nli 46.2 46.3 35.4
Average 71.2 (4.3$\uparrow$) 71.5 (4.6$\uparrow$) 66.9
We compare three versions of the LLAMA2-7B-chat-hf model: the original non-fine-tuned model, a version fine-tuned using SFT, and another fine-tuned with PPO. The 95% confidence intervals (CI) reported here are defined by the 2.5th and 97.5th percentiles of the bootstrap distribution. Using a consistent prompt across models, we find that the PPO-fine-tuned model achieves an accuracy of 39.35% (95% CI: 38.39, 40.29), significantly outperforming the original model, which achieves 27% accuracy (95% CI: 19.00, 36.03). Conversely, the SFT-fine-tuned model demonstrates extremely poor performance, achieving less than 1% accuracy.
Method Accuracy 95% CI
Original 27.00 (19.00, 36.03)
SFT 0.00961 (0.00, 0.03)
PPO 39.35 (38.39, 40.29)
Qualitative analysis of sampled outputs reveals that the PPO-fine-tuned model reliably adheres to the instruction format and generates detailed reasoning to support its predictions. In contrast, the SFT-fine-tuned model often fails to adapt its responses to the required format, demonstrating limited generalization capabilities. PPO fine-tuning maintains proximity to the original model distribution via a clipping mechanism, thus preserving and enhancing the model’s intrinsic instruction-following capabilities. In contrast, SFT fine-tuning appears to narrow the model’s learned distribution to task-specific training data, negatively impacting its original instruction-following proficiency. We also evaluate the impact of fine-tuning on language modeling ability in Appendix E.
5.6 Performance Comparison Across Different LLMs
@@
To assess the consistency of our findings across different models, we evaluated Qwen2.5-7B-Instruct and MPT-7B-chat alongside LLAMA2-7B-chat-hf on the STS-B dataset from the GLUE benchmark and the COPA dataset from the SuperGLUE benchmark. The results confirm that PPO-based fine-tuning consistently outperforms the BERT-large model, as well as the zero-shot and few-shot prompting baselines for all LLMs, highlighting its effectiveness across different LLMs. Additionally, the effect of few-shot prompting on COPA performance varies across different LLMs, indicating that different LLMs have varying capabilities to process and follow long-context instructions, which results in variable performance outcomes.
6 Conclusion
Limitations
This work takes an initial step toward framing NLU as a reinforcement learning problem for LLMs under 14B parameters. While our long-term goal is to reduce reliance on curated datasets by leveraging richer, task-specific reward models, we currently adopt a simple binary reward signal based on exact label matching. This design enables a controlled evaluation of PPO as an effective adaptation strategy, showing consistent gains over prompting and supervised fine-tuning. Although we present a preliminary exploration of model-driven reward functions in Appendix G, further research is needed to develop robust and generalizable reward signals that can support more complex or weakly supervised tasks without requiring extensive manual annotation. Overall, our findings suggest that casting nuanced tasks as reinforcement learning problems—through the design of appropriate environments and reward functions—offers a scalable and flexible alternative to standard fine-tuning, particularly when the model is already well-initialized through pretraining.
Appendix A Hyperparameter Settings
@@
Financial PhraseBank (96.2, 98.1) (96.9, 98.5) (96.6, 98.4)
Labelled Financial News (66.1, 74.6) (69.0, 76.6) (63.2, 72.2)
Mental Health (66.6, 67.7) (66.0, 67.1) (59.3, 60.5)
Emotion (77.4, 78.6) (75.8, 77.0) (77.0, 78.2)
@@
Babi-nli (64.3, 71.5) (65.1, 73.0) (58.8, 67.6)
SIGA-nli (39.0, 53.9) (40.6, 53.7) (28.5, 42.2)
Appendix B Reward Curve for PPO Fine-Tuning
Appendix C Results on SuperGLUE Benchmark
Models BoolQ CB COPA MultiRC ReCoRD RTE
BERT-large 77.4 75.7/83.6 70.6 70.0/24.0 72.0/71.3 71.6
BERT-large++ 79.0 84.7/90.4 73.8 70.0/24.1 72.0/71.3 79.0
LLAMA2-7B-chat-hf
Zero-shot prompting 75.8 26.4/43.6 57.0 51.9/20.3 27.0/26.2 59.2
Few-shot prompting 80.2 49.8/66.0 73.4 46.6/15.4 36.3/35.3 72.9
PPO-ST 85.9 74.7/88.0 88.6 82.5/50.0 70.6/69.9 84.3
Models WiC WSC AXb AXg Average
BERT-large 69.5 64.3 23.0 97.8/51.7 69.0
BERT-large++ 69.5 64.3 38.0 99.4/51.4 71.5
LLAMA2-7B-chat-hf
Zero-shot prompting 54.4 52.1 9.1 64.0/55.1 49.5
Few-shot prompting 54.4 62.3 9.1 64.0/55.1 54.9
PPO-ST 72.1 78.1 52.7 91.0/79.8 78.3
We fine-tuned the LLAMA2-7B-chat-hf model using PPO on the SuperGLUE dataset and compared its performance against several baselines, including BERT-large, BERT-large++, and zero-shot and few-shot prompting of LLAMA2-7B-chat-hf. The term “BERT++” refers to a BERT model fine-tuned using the supplementary training on intermediate labeled-data tasks (STILTs) approach (Phang et al., 2018), where the model is first fine-tuned on related transfer tasks before being fine-tuned on SuperGLUE tasks. For example, MNLI from the GLUE benchmark(Wang et al., 2019) is used as an intermediate task for CB, RTE, and BoolQ(Wang et al., 2020). In contrast, our experiments with LLM did not use this method. Our models were only fine-tuned on the datasets included in the SuperGLUE benchmark.
It is worth noting that on MultiRC, few-shot prompting performs slightly worse than zero-shot prompting. This may be because MultiRC involves long input contexts, and incorporating multiple examples in a few-shot prompt can cause the total input length to approach or exceed the LLMs maximum context window. Even in the one-shot setting, providing an excessively long context can dilute the model’s attention, potentially leading to reduced performance.
Appendix D Evaluation on Reading Comprehension Tasks
We evaluate LLAMA2-7B-chat-hf on the SQuAD reading comprehension task, where the objective is to select a passage from a given context that best answers a question. Two training strategies are compared: Supervised Fine-Tuning (SFT), which directly uses the ground-truth answer as the training label, and Proximal Policy Optimization (PPO), which leverages reward functions based on Exact Match (EM) and F1 Score. EM metric is computed by comparing a normalized prediction against the normalized ground truth (with normalization involving lowercasing and punctuation removal); a perfect match yields an EM score of 1, otherwise 0. F1 score measures word-level overlap, balancing how many predicted words are correct (precision) and how many ground-truth words are included (recall).
Method EM F1
Original 7.66 32.27
SFT 59.17 76.48
PPO 65.74 81.82
Appendix E Impact of Fine-Tuning on Language Modeling Ability

### v5

@@ Abstract
Instruction-fine-tuned large language models (LLMs) under 14B parameters continue to underperform on natural language understanding (NLU) tasks, often trailing smaller models like BERT-base on benchmarks such as GLUE and SuperGLUE. Motivated by the success of reinforcement learning in reasoning tasks (e.g., DeepSeek), we explore Proximal Policy Optimization (PPO) as a framework to improve the NLU capabilities of LLMs. We frame NLU as a reinforcement learning environment, treating token generation as a sequence of actions and optimizing for reward signals based on alignment with ground-truth labels. PPO consistently outperforms supervised fine-tuning, yielding an average improvement of 6.3 points on GLUE, and surpasses zero-shot and few-shot prompting by 38.7 and 26.1 points, respectively. Notably, PPO-tuned models outperform GPT-4o by over 4% on average across sentiment and natural language inference tasks, including gains of 7.3% on the Mental Health dataset and 10.9% on SIGA-nli. This work highlights a promising direction for adapting LLMs to new tasks by reframing them as reinforcement learning problems, enabling learning through simple end-task rewards rather than extensive data curation. Our code is available at https://github.com/coder-qicao/RL4GLUE.
@@
Sai Ashish Somayajula**footnotemark: *  Bokai Hu [^1] Equal Contribution   Qi Cao   Xin Pan   Pengtao Xie UC San Diego {ssomayaj, boh001, q9cao, p1xie}@ucsd.edu
@@ 1 Introduction
Figure 1: Despite extensive pre-training and instruction tuning, zero-shot and few-shot prompting of models like LLAMA2-7B-chat-hf and LLAMA2-13B-chat-hf continues to underperform on NLU tasks, often falling short of smaller encoder-based models like BERT. To address this performance gap, we recast the task of NLU adaptation as an RL problem, fine-tuning the model using PPO. The input consists of a prompt and a query. In the standard setting (top path), the base model, with zero-shot/few-shot prompting, struggles to generate correct answers. Our approach selectively updates lightweight LoRA layers, using PPO to optimize a reward signal based on task correctness. This optimization encourages performance gains while constraining deviation from the base policy. We find that PPO fine-tuning substantially outperforms standard SFT.
Large language models (LLMs) (Radford et al., 2019; Brown, 2020; Touvron et al., 2023b) have revolutionized natural language processing (NLP) with their powerful text generation capabilities (Radford, 2018). Pretrained on large-scale unlabeled corpora, LLMs can generate coherent and contextually relevant content. Using prompt-based strategies like zero-shot and few-shot prompting (Brown, 2020), these models can address a wide range of downstream tasks without task-specific fine-tuning. However, when applied to instruction-fine-tuned LLMs under 14B parameters  [^2] The 14B cutoff reflects our computational constraints; LLAMA-2-13B-chat-hf is the largest model we were able to evaluate.—such as LLAMA2-7B-chat-hf—these methods often underperform on natural language understanding (NLU) tasks compared to encoder-only models like BERT (Devlin, 2018), which consistently excel on benchmarks such as GLUE (Wang et al., 2019) and SuperGLUE (Wang et al., 2020). For instance, our evaluation of LLAMA2-7B-chat-hf shows that zero-shot prompting with task-specific prompts yields an average performance of 46.1 across all GLUE datasets, while few-shot prompting improves performance to 58.7—both significantly trailing BERT-base’s 79.6, as shown in Table 1.
@@
While standard fine-tuning (SFT) is commonly used to adapt LLMs to downstream tasks, we find it insufficient for NLU—often underperforming even smaller encoder-only models like BERT-base. In contrast, we use PPO to enhance LLM performance by framing NLU as a reinforcement learning problem, as shown in Fig 1. The sequence of input tokens up to timestep t1t-1 represents the state $s_{t}$, and the token generated at timestep $t$ is treated as the action $a_{t}$. After generating the full response, a heuristic extracts the predicted answer, which is compared to the ground-truth label to assign a scalar reward $R$. PPO then updates the model to maximize this reward, enabling direct optimization for task-specific objectives. Empirically, PPO-based fine-tuning of LLAMA2-7B-chat-hf improves GLUE performance by 6.3 points over SFT, surpasses zero and few-shot prompting by 38.7 and 26.1 points, respectively, and even outperforms GPT-4o by over 4% on sentiment and inference tasks—achieving gains of 7.3% on the Mental Health dataset and 10.8% on SIGA-nli. These results demonstrate the effectiveness of reinforcement learning in aligning LLMs under 14B parameters with NLU objectives.
@@
Pre-trained LLMs possess broad linguistic knowledge, spanning syntactic and semantic structures, acquired from large-scale text corpora. We show that reinforcement learning, specifically PPO, can refine this general understanding to better align with task-specific NLU objectives. Similar patterns are observed in DeepSeek, where chain-of-thought pre-training enhances reasoning capabilities, and subsequent RL-based fine-tuning further improves performance. Building on this insight, our findings suggest a promising direction: adapting LLMs to new tasks by formulating them as reinforcement learning problems. When models are sufficiently pre-trained, task alignment may be achieved without additional labeled data—requiring only a well-defined reward function over the outputs. PPO can then optimize the model toward high-reward behaviors. This approach offers a scalable, label-efficient alternative to conventional supervised fine-tuning through reward-driven adaptation.
@@ 2 Method
To enhance the performance of LLMs on NLU tasks, we adopt two distinct fine-tuning methods. The first approach involves supervised fine-tuning, where the input consists of a concatenation of the task-specific prompt, query and the ground truth answer, with the model optimized using the next-token prediction objective. The second approach utilizes PPO, framing response generation as a reinforcement learning problem. In this setup, the sequence of input tokens until timestep t1t-1 represents the state $s_{t}$, and each token generated at timestep $t$ is treated as an action $a_{t}$. After generating the entire sequence, a heuristic-based process extracts the final answer from this generated sequence, and is compared to the ground truth. PPO is then employed to optimize the model by maximizing the cumulative reward derived from this comparison. To reduce computational complexity, we fine-tune LoRA layers instead of the full model. We refer the readers to Appendix A and B for preliminaries on PPO.
@@ 2.3 Proximal Policy Optimization for LLM Fine-tuning on NLU Tasks
We utilize PPO to fine-tune the LLM on NLU tasks, following the training protocol outlined in Appendix B. The reward function is specifically designed for each NLU task. In this work, we use a simple reward function, where a reward is assigned at the end of the generation based on alignment with the ground truth labels. We use regular expression matching to extract answers from the LLMs outputs by first locating the text within the ‘<Judgement></Judgement>’ tags. Depending on the task, we then search for task-specific keywords (such as “yes”, “no”, “acceptable”, or “not acceptable”) to identify the answer. These extracted answers are compared with the ground truth to determine the appropriate rewards.
For instance, CoLA, a classification task, answers are categorized as acceptable, unacceptable, or exceptional (incorrect format). For STS-B, a regression task, the extracted answer is a floating-point number between 0 and 5. Reward per generation for classification tasks is given by $R=\mathds{1}(\hat{y}==y_{i})$, where $\hat{y}$ is model’s prediction and $y$ is ground truth. For STS-B, a regression task, the reward for each generation is defined as $R=2.5-|\hat{y}_{i}-y_{i}|$, where $\hat{y}_{i}$ is the predicted score and $y_{i}$ is the ground truth. Since both scores lie in $[0,5]$, their absolute difference ranges from 0 to 5. Subtracting this from 2.5 centers the reward around zero and bounds it within [2.5, 2.5][-2.5,\,2.5], with the maximum reward achieved when $\hat{y}_{i}=y_{i}$. Incorrectly formatted responses are penalized with a value of -1 for classification tasks and -2.5 for regression tasks, representing the largest penalties applied in each case.
@@ 3.1 Experimental Setup
We trained and evaluated our models on the GLUE (Wang et al., 2019) and SuperGLUE (Wang et al., 2020) benchmarks. All experiments were conducted using instruction-tuned LLAMA2-7B models (Touvron et al., 2023a) [^3] https://huggingface.co/daryl149/llama-2-7b-chat-hf. We perform both single task and multi-task fine-tuning: 1) Single-task fine-tuning: For each subtask within GLUE and SuperGLUE, a separate task-specific LoRA module was trained independently. 2) Multi-task fine-tuning: In the multi-task setting, datasets from different subtasks within each benchmark were combined, and a single LoRA module was trained to handle all tasks simultaneously. Please refer to Appendix C for detailed hyperparameter settings.
@@ 3.3 Results on GLUE Benchmark
Models MNLI-m MNLI-$m$ QQP QNLI SST-2 CoLA
@@
To assess the consistency of our findings across different models, we evaluated Qwen2.5-7B-Instruct and MPT-7B-chat alongside LLAMA2-7B-chat-hf on the STS-B dataset from the GLUE benchmark and the COPA dataset from the SuperGLUE benchmark. The results are summarized in Table 2. The results confirm that PPO-based fine-tuning consistently outperforms the BERT-large model, as well as the zero-shot and few-shot prompting baselines for all LLMs, highlighting its effectiveness across different LLMs.
3.4 Comparison of RL Algorithms: PPO vs. GRPO
Algorithm SST-2 MRPC RTE CoLA QNLI Avg. Per-Step Runtime (s)
SFT 73.8 85.8 80.4 50.7 93.6 76.9 4.124
PPO 96.4 89.4 84.3 59.9 93.2 84.6 4.299
GRPO 96.7 91.2 88.5 55.2 93.1 84.9 5.155
Table 3: Comparison of SFT, PPO, and GRPO fine-tuning methods on the LLAMA2-7B-chat-hf model across five GLUE benchmark tasks (SST-2, MRPC, RTE, CoLA, QNLI), along with per-step runtime.
Our objective is to improve the natural language understanding capabilities of the base (policy) model through RL fine-tuning. In this context, we compare two RL approaches: PPO and Group Relative Policy Optimization (GRPO)  (Shao et al., 2024). PPO requires updating a separate critic model to compute value functions, which introduces (modest) additional memory constraints. On the other hand, GRPO was designed to bypass the critic model entirely. Instead, GRPO samples multiple trajectories per prompt and computes each trajectory’s advantage by comparing its reward to the batch’s average (and standard deviation). This method not only simplifies the architecture but also reduces memory usage.
As detailed in Table 3, both PPO and GRPO deliver notable performance improvements over SFT. Notably, PPO only incurs about a 4% increase in per-step runtime compared to SFT. However, GRPO’s need to generate multiple responses per sample results in a higher runtime, despite its memory efficiency. Overall, our analysis highlights the trade-offs between these RL algorithms: PPO offers efficient runtime with the cost of additional overhead from the critic model, while GRPO reduces memory usage at the expense of increased sampling time.
@@ 3.5 Evaluating Zero-Shot Generalization of PPO Fine-Tuned Models and Comparison with GPT-4o
Financial PhraseBank 97.2 97.7 97.5
Labelled Financial News 70.2 72.3 67.8
Mental Health 67.2 66.6 59.9
Emotion 78.0 76.4 77.6
@@
Babi-nli 68.3 69.4 63.2
SIGA-nli 46.2 46.3 35.4
[…]
We compare zero-shot prompting of three versions of the LLAMA2-7B-chat-hf model: the original non-fine-tuned model, a version fine-tuned using SFT, and another fine-tuned with PPO. The 95% confidence intervals (CI) reported here are defined by the 2.5th and 97.5th percentiles of the bootstrap distribution. Using a consistent prompt template across models, we find that the PPO-fine-tuned model achieves an accuracy of 39.35% (95% CI: 38.39, 40.29), significantly outperforming the original model, which achieves 27% accuracy (95% CI: 19.00, 36.03). Conversely, the SFT-fine-tuned model demonstrates extremely poor performance, achieving less than 1% accuracy.
Method Accuracy 95% CI
Non-fine-tuned 27.00 (19.00, 36.03)
SFT 0.00961 (0.00, 0.03)
PPO 39.35 (38.39, 40.29)
Qualitative analysis of sampled outputs reveals that the PPO-fine-tuned model reliably adheres to the instruction format and generates detailed reasoning to support its predictions. In contrast, the SFT-fine-tuned model often fails to adapt its responses to the required format, demonstrating limited generalization capabilities. PPO fine-tuning maintains proximity to the original model distribution through its clipping mechanism and the KL-divergence minimization term in the objective function (part of the TRL library), thereby preserving and enhancing the model’s intrinsic instruction-following capabilities. In contrast, SFT fine-tuning appears to narrow the model’s learned distribution to task-specific training data, negatively impacting its original instruction-following proficiency.
3.7 Impact of Fine-Tuning on Language Modeling Ability